OpenStudy (anonymous):

Simplify. Square root 9 / 6 square root 9

OpenStudy (xapproachesinfinity):

root9=3 the rest is easy

OpenStudy (anonymous):

I not sure. Can you explain?

OpenStudy (xapproachesinfinity):

square root 9=3 so it becomes 3/(6 times 3) \[\frac{3}{6\times 3}\]

OpenStudy (xapproachesinfinity):

do the rest

OpenStudy (anonymous):

@zapproachesinfinity

OpenStudy (anonymous):

I got 1/6 but the answer possibilities are: 9 -2/3 9 -1/3 9 1/2 9 1/3

OpenStudy (xapproachesinfinity):

well then you wrote the question incorrectly is it \(\huge \frac{\sqrt{9}}{6\sqrt{9}}\)

OpenStudy (anonymous):

Here

OpenStudy (anonymous):

Sorry... Can you help me?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

here \(\huge \frac{\sqrt{9}}{\sqrt[6]{9}}=\frac{3}{3^{\frac{2}{6}}}=\frac{3}{3^{1/3}}\)

OpenStudy (xapproachesinfinity):

yes i'm here just working on writing the solution

OpenStudy (anonymous):

So the answer is d?

OpenStudy (xapproachesinfinity):

no im not finished yet! sorry the the web site is slow

OpenStudy (xapproachesinfinity):

\[\huge 3^{1-\frac{1}{3}}=3^{\frac{2}{3}=(3^2)^{1/3}}=9^{1/3}\]

OpenStudy (xapproachesinfinity):

@Nicole2525

OpenStudy (anonymous):

Thank you... So the answer or d?

OpenStudy (anonymous):

Is*